OpenStudy (anonymous):

how do i put that in slope intercept form (4,7; m = - 3/4

OpenStudy (igreen):

Put it in point-slope form first. \(\sf y - y_1 = m(x - x_1)\) Where \(\sf y_1\) is the y-value of the point, \(\sf x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(\sf y - 7 = -\dfrac{3}{4}(x - 4)\) Distribute -3/4

OpenStudy (anonymous):

Ok so do you know what the slope and the y-intersept?

OpenStudy (anonymous):

are